OpenStudy (anonymous):

(x^2-1)^1/2 - (x^2-1)^-1/2(2-x^2) all divided by x^2-1

hero (hero):

It's all you Jim

jimthompson5910 (jim_thompson5910):

Is the expression \[\Large \frac{(x^2-1)^{\frac{1}{2}} - \frac{(x^2-1)^{-\frac{1}{2}}}{(2-x^2)}}{x^2-1}\] or is it \[\Large \frac{\frac{(x^2-1)^{\frac{1}{2}} - (x^2-1)^{-\frac{1}{2}}}{(2-x^2)}}{x^2-1}\] ???

jimthompson5910 (jim_thompson5910):

or is it \[\large \frac{(x^2-1)^{\frac{1}{2}} - (x^2-1)^{-\frac{1}{2}}(2-x^2)}{x^2-1}\]

OpenStudy (anonymous):

second equation and i need to solve for x

jimthompson5910 (jim_thompson5910):

You can't solve for x here because this isn't an equation. You can only simplify. Make sure you've copied the entire problem correctly.

OpenStudy (anonymous):

it says the answer is \[\pm \sqrt{3/2}\] and it says solve for x and and in parentheses it in says put in simplest form.

jimthompson5910 (jim_thompson5910):

Ok,so that means that this is an equation, but I don't see any equal signs. Can you repost the actual equation?

OpenStudy (anonymous):

\[(x ^{2}-1)^{1/2}-(x ^{2}-1)^{-1/2}(2-x ^{2}) \div (x ^{2}-1)= 0\]

jimthompson5910 (jim_thompson5910):

ah ok, thx

jimthompson5910 (jim_thompson5910):

\[\large (x^{2}-1)^{1/2}-(x^{2}-1)^{-1/2}(2-x^{2}) \div (x^{2}-1)= 0\] \[\large (x^{2}-1)^{1/2}-(x^{2}-1)^{-1/2}(2-x^{2})= 0*(x^{2}-1)\] \[\large (x^{2}-1)^{1/2}-\frac{1}{(x^{2}-1)^{1/2}}(2-x^{2})= 0\] \[\large \sqrt{x^{2}-1}-\frac{1}{\sqrt{x^{2}-1}}(2-x^{2})= 0\] \[\large x^{2}-1-(2-x^{2})= 0\] \[\large x^{2}-1-2+x^{2}= 0\] \[\large 2x^{2}-3= 0\] \[\large 2x^{2}= 3\] \[\large x^{2}= \frac{3}{2}\] \[\large x= \pm\sqrt{\frac{3}{2}}\]

OpenStudy (anonymous):

did the two radicals cancel out?

jimthompson5910 (jim_thompson5910):

I multiplied EVERY term by \[\sqrt{x^2-1}\] to clear out the fractions. In doing that, the first term became \[\sqrt{x^2-1}*\sqrt{x^2-1}=(\sqrt{x^2-1})^2=x^2-1\] For the second term, they simply canceled.

OpenStudy (anonymous):

oh ok thank you so much